# CdTe and HgTe doped with V, Cr, and Mn – prospects for the quantum anomalous Hall effect

### v1

@@ IV Local distortions produced by the Jahn-Teller effect
Lattice distortions produced by the Jahn-Teller (JT) effect in zinc-blende II-IV semiconductors are not straightforward. The impurities in ZnSeS were shown to produce octahedral rotations and changes of bond-angles Sarkar et al. (2022). Both Cr- and V-doping induce moderate tetrahedral distortions in CdTe and smaller tetrahedral distortions in HgTe. The main effect is a change in the Te-M-Te bond angle. In regular tetrahedra, we have four angles equal to 109.5 degrees. Among these four angles, two remain almost unchanged in the case of doping, one of these angles becomes smaller and another becomes larger. In Fig. 4 we show the crystal structure and the angles that change with the JT distortion.
| CdCrTe | CdVTe | HgCrTe | HgVTe | |
| Bond length (Å) | ||||
| M-Te | 2.8019 | 2.7622 | 2.8028 | 2.7282 |
| M-Te | 2.8040 | 2.7520 | 2.8057 | 2.7266 |
| M-Te | 2.7733 | 2.7600 | 2.7589 | 2.7271 |
| Angles (deg) | ||||
| 102.80 | 112.60 | 101.13 | 109.73 | |
| 115.72 | 107.64 | 117.07 | 109.25 | |
| PV(Å3) | 11.053 | 10.739 | 10.932 | 10.409 |
| BAV (deg) | 43.054 | 2.794 | 69.191 | 0.025 |
| DI (bond length) | 0.0040 | 0.0017 | 0.0062 | 0.0002 |
| (meV) | 34.696 | 4.400 | 45.05 | 2 20.094 |
When we dope the systems with Cr, the angle becomes larger and becomes smaller, the opposite happens when we dope with V. This situation is different from what happens in other systems Forte et al. (2018); Ni et al. (2020), where the two angles and become larger or smaller together. This allows a breaking of the degeneracy of the energy levels compatible with the cubic symmetry of the system. Two larger angles or two smaller angles will be more efficient in producing the JT splitting, and they will create a local tetragonal atom arrangement. In Table 1, we describe the local distortions in different systems and the gain in the total energy resulting from the JT distortion. These JT distortions help in stabilizing the localized character of states derived from TM levels.
@@ V Exchange couplings between transition metal spins
In this section, we present computational results on exchange couplings between TM spins employing . In Table 2, we report energy differences between antiferromagnetic (AFM) and ferromagnetic (FM) configurations of TM spins in CdTe and HgTe doped with V, Cr, and Mn without taking the Jahn-Teller distortion into account. As seen, FM couplings prevail for the V case, whereas the interaction is AFM for Cr- and Mn-doped compounds. Once the Jahn-Teller distortion is taken into account, FM couplings show up in the Cr-compounds, too, as shown in Table 3. A comparison of data for presented here to results obtained for and 0.50 summarized in Appendix, demonstrate a strong sensitivity of the coupling strength to the employed theoretical framework. This sensitivity confirms the presence of a delicate balance between FM and AFM contributions to the TM coupling in the case of early transition metals in II-VI compounds Śliwa and Dietl (2023).
@@
| CdVTe | -22.3 | 78.3 | -3.4 |
| CdCrTe | -57.9 | -22.6 | -1.6 |
| CdMnTe | -16.2 | -16.1 | -0.3 |
| HgVTe | 35.5 | 41.2 | -2.7 |
| HgCrTe | -313.4 | -150.1 | -4.0 |
| HgMnTe | -17.0 | -16.9 | -1.6 |
| compound | ||
| CdCrTe NO JT | -57.9 | -22.6 |
| CdCrTe JT/2 | 10.4 | 4.8 |
| CdCrTe JT | 5.9 | 34.5 |
| HgCrTe NO JT | -313.4 | -150.1 |
| HgCrTe JT/2 | -100.4 | -154.4 |
| HgCrTe JT | -54.8 | -231.0 |
VI Conclusions
According to theoretical studies presented in the present and the companion paper Śliwa and Dietl (2023), the predicted magnitude and often sign of magnetic coupling between Cr ions in II-VI compounds depends on the approach employed. This fact reflects a competition between ferromagnetic superexchange and mostly antiferromagnetic Bloembergen-Rowland and two-electron terms Śliwa and Dietl (2023). Furthermore, according to results presented in Appendix in Fig. 6, pinning of the Fermi level below the top of the valence band by Cr ions in (Hg,Cr)Te cannot be entirely excluded. Ferromagnetic interactions appear more robust in V-doped compounds, though the donor character of V-impurities may show up in II-VI tellurides.
The ensemble of our results indicates that the most promising strategy for obtaining a QAHE device of HgTe-based systems is to employ (Cd,Hg,Cr)Te barriers that should be ferromagnetic already above 4.2 K, and in which - exchange splitting of bands should lead to splitting of QW states. That splitting can be further enhanced by a few percent Cr doping of the HgTe QW. However, the observation of the QAHE may require tilting of magnetization away from in-plane or perpendicular orientation Śliwa and Dietl (2023). Doping by V is expected to lead to a more robust ferromagnetic ground state but it could be that the resulting electron doping may preclude the shifting of the Fermi level to the topological gap. In any case, a close energetic proximity of states and the Fermi level opens doors for a new physics in both (Hg,V)Te and (Hg,Cr)Te, not encountered in the case of (Hg,Mn)Te.
@@ Acknowledgments
The work is supported by the Foundation for Polish Science through the International Research Agendas program co-financed by the European Union within the Smart Growth Operational Programme (Grant No. MAB/2017/1). We acknowledge the access to the computing facilities of the Interdisciplinary Center of Modeling at the University of Warsaw, Grant G84-0, GB84-1, and GB84-7. We acknowledge the CINECA award under the ISCRA initiative IsC85 ”TOPMOST” and IsC93 ”RATIO” grant, for the availability of high-performance computing resources and support. We acknowledge the access to the computing facilities of the Poznan Supercomputing and Networking Center Grant No. 609.
@@ Appendix A Transition metal doped compounds with 25% and 50% of the exact exchange
In this Appendix, we present band structures and the magnetic properties for TM-doped CdTe and HgTe obtained employing the hybrid functional with and of the exact exchange. The band structure results displayed in Figs. 5, 6, and 7 substantiate the use of in the main body of the text. In particular, in the case of HgMnTe, the experimental value of the band gap eV Furdyna (1988) is close to the computational data for (Fig. 3) but is too small and too large for and 0.5, respectively, as shown in Figs. 6 and 7.
@@
| CdVTe | -38.1 | 22.8 | 0.6 |
| CdCrTe | -10.5* | -16.8* | -4.8 |
| CdMnTe | -21.9 | -21.8 | -0.5 |
| HgVTe | 47.6 | 51.5 | -1.6 |
| HgCrTe | -77.9 | -35.9 | -5.1 |
| HgMnTe | -22.3 | -22.2 | -2.0 |
@@
| CdCrTe NO JT | -218.7 | -160.3 |
| CdCrTe JT/2 | -62.3 | 24.1 |
| HgCrTe NO JT | 74.0 | 7.8 |
| HgCrTe JT/2 | 88.6 | -22.7 |

### v2

@@ IV Local distortions produced by the Jahn-Teller effect
While the long-range cooperative Jahn-Teller (JT) effect in homogeneous bulk systems can be described with a finite number of phonon modes as in molecules, the description of the JT effect in solids with JT-active defects involves an infinite number of phonon modes. A particularly tricky case is represented by the case of JT-active impurities in tetrahedra environments like CdTe and HgTe which are analyzed in this paper. Recently, Cruz et al.Cruz et al. (2024) presented a recipe to tackle this problem in the case of topologically trivial CdSe which is similar to CdTe, while we are not aware of attempting to solve the same problem in topologically non-trivial systems like HgTe. Therefore, the lattice distortions produced by the JT effect in zinc-blende II-IV semiconductors are not straightforward and the effect can be relevant. For instance, the impurities in zinc-blende ZnSexS1-x were shown to produce octahedral rotations and changes of bond-angles Sarkar et al. (2022). The JT lifted the degeneracy of the t2g and eg manifold for both spins in both Cr and V doped systems, this was established through the check of the electronic levels at the point. The lifting of the degeneracy of the manifold is necessary to observe an insulating phase. The main effect is a change in the Te-M-Te bond angle. In regular tetrahedra, we have four angles equal to 109.5 degrees. Among these four angles, two remain almost unchanged in the case of doping, one of these angles becomes smaller and another becomes larger. In Fig. 5 we show the crystal structure and the angles that change with the JT distortion. The bond angle variance (BAV) is much larger for the V-doping than for the Cr-doping. While the eg manifold for the majority spin is full, the V-doping consists of 1 electron in the t2g manifold, while the Cr-doping consists of 2 electrons in the t2g manifold. We assume that the BAV is larger for the case of a single electron in the manifold as happens for the d4 in an octahedral crystal field.
When we dope the systems with Cr, the angle becomes larger and becomes smaller, the opposite happens when we dope with V. This situation is different from what happens in other systems Forte et al. (2018); Ni et al. (2020), where the two angles and become larger or smaller together. The JT allows the breaking of the degeneracy of the energy levels compatible with the cubic symmetry of the system. Two larger angles or two smaller angles will be more efficient in producing the JT splitting, and they will create a local tetragonal atom arrangement. In Table 1, we describe the local distortions in different systems and the gain in the total energy resulting from the JT distortion. These JT distortions help in stabilizing the localized character of states derived from TM levels.
@@ V Exchange couplings between transition metal spins
In this Section, we present computational results on exchange couplings between TM spins employing . In Table 2, we report energy differences between antiferromagnetic (AFM) and ferromagnetic (FM) configurations of TM spins in CdTe and HgTe doped with V, Cr, and Mn without taking the Jahn-Teller distortion into account. As seen, FM couplings prevail for the V case, whereas the interaction is AFM for Cr- and Mn-doped compounds. Once the Jahn-Teller distortion is taken into account, FM couplings show up in the Cr-compounds, too, as shown in Table 3. Regarding the V-doped systems, the ground state is still FM if we include the Jahn-Teller distortions, as shown in Table 4. A comparison of data for presented here to results obtained for and 0.50 summarized in Appendix, demonstrate a strong sensitivity of the coupling strength to the employed theoretical framework. This sensitivity confirms the presence of a delicate balance between FM and AFM contributions to the TM coupling in the case of early transition metals in II-VI compounds Śliwa and Dietl (2023).
@@
| Cd1-xVxTe | -22.3 | 78.3 | -3.4 |
| Cd1-xCrxTe | -57.9 | -22.6 | -1.6 |
| Cd1-xMnxTe | -16.2 | -16.1 | -0.3 |
| Hg1-xVxTe | 35.5 | 41.2 | -2.7 |
| Hg1-xCrxTe | -313.4 | -150.1 | -4.0 |
| Hg1-xMnxTe | -17.0 | -16.9 | -1.6 |
| compound | |||
| Cd1-xCrxTe NO JT | -57.9 | -22.6 | -1.6 |
| Cd1-xCrxTe JT/2 | 10.4 | 4.8 | -1.6 |
| Cd1-xCrxTe JT | 5.9 | 34.5 | -1.7 |
| Hg1-xCrxTe NO JT | -313.4 | -150.1 | -4.0 |
| Hg1-xCrxTe JT/2 | -100.4 | -154.4 | -4.6 |
| Hg1-xCrxTe JT | -54.8 | -231.0 | -5.6 |
| compound | |||
| Cd1-xVxTe NO JT | -22.3 | 78.3 | -3.4 |
| Cd1-xVxTe JT | 28.3 | 28.4 | -3.1 |
| Hg1-xVxTe NO JT | 35.5 | 41.2 | -2.7 |
| Hg1-xVxTe JT | 0.2 | 30.5 | -7.5 |
VI Implications to the anomalous Hall effect
Quantum wells of HgTe constitute a rare example of systems showing the quantum spin Hall effect. Our studies presented here and in the companion paper Śliwa and Dietl (2023) were designed to address the question about the suitability of HgTe QWs and related systems for the demonstration of the quantum anomalous Hall effect (QAHE) with characteristics opening a door or applications in the quantum metrology. As mentioned in Introduction, this question is timely considering the fact that Hall resistance quantization in (Bi,Sb,Cr,V) layers is perturbed above 50 mK by hopping conductivity Kawamura et al. (2017) involving numerous in-gap states brought about by antisites and other defects Zhang et al. (2012). As already noted by Liu et al. Liu et al. (2008), the appearance of the chiral edge states requires the presence of: (1) localized spin polarization that breaks the time-reversal symmetry and generates giant exchange spin splitting of bands; (2) a gaped inverted band structure in one spin channel (-like states above -like states), and the gaped normal band arrangement in the other spin channel – to achieve this situation a proper ratio between - and - exchange integrals is necessary; (3) the Fermi level in the gap region. That pioneering theory of the QAHE considered (Hg,Mn)Te QW as a system in which that phenomenon could be observed Liu et al. (2008), albeit in the presence of an external magnetic field, as interactions between Mn spins are antiferromagnetic according to our and previous studies Mycielski et al. (1984); Śliwa et al. (2021). However, recent experimental findings have indicated that the QHE dominates in (Hg,Mn)Te and becomes visible already in 50 mT Shamim et al. (2020), the fact linked to the resonant character of acceptor states and the formation of bound magnetic polarons Dietl (2023a, b). Under these conditions, (Hg,Mn)Te looks as a promising QHE resistance standard working in surprisingly low magnetic fields, though its performance above 50 mK is to be verified experimentally. Our first-principles results for (Hg,V)Te point to the presence of a FM coupling between V spins, an encouraging finding from the QAHE perspective. At the same time, however, our data reveal that the V impurities form a resonant donor state with the conduction band. This would mean that V doping will introduce electrons that would preclude a shift of the Fermi level toward the topological gap by a gate voltage. In the case of (Hg,Cr)Te, our ab initio studies indicate that the Cr donor level is degenerate with the valence band and, therefore, Cr acts as an isoelectronic impurity, similarly to Mn. Accordingly, it should be possible to shift the Fermi level towards the topological gap either by a suitable co-doping with electrically active impurities or by gating. At the same time, our first principles data substantiate two conclusions of the tight-binding computations Śliwa and Dietl (2023): (i) The exchange coupling between Cr spins is on the borderline between the FM case of V ions and AFM interactions between Mn ions. This opens a door for manipulation of magnetism by an electric field, strain, or pressure but suggests that the application of an external magnetic field may be necessary to polarize Cr spins. (ii) A close proximity of the Cr donor level to the top of the valence band results in a large - hybridization and, thus, in a large - exchange interaction between band carriers and localized spins, which we see as large splittings of and bands in Figs. 3 and 4. It has been shown in the companion paper Śliwa and Dietl (2023), exploiting the methodology developed earlier for HgTe and (Hg,Mn)Te QWs Novik et al. (2005); Dietl (2023b), that a large ratio between the - and - exchange integrals (denoted as and in the literature) precludes the appearance of the QAHE for the magnetization orientation along the growth direction [001] in (Hg,Cr)Te/(Cd,Hg)Te QWs. However, the spin-orbit interaction reduces the - splitting for a tilted magnetization direction. A direct calculation for the [111] magnetization direction, 3 of Hg-substitutional Cr ions in HgTe QW with a thickness of 6 nm, has demonstrated the presence of the gaped inverted band structure in one spin channel and the normal gaped band arrangement for the other spin directionŚliwa and Dietl (2023). We conclude, therefore, that (Hg,Cr)Te QWs are a perspective system for the observation of the QAHE. Furthermore, the determined magnitude of the topological energy gap of the order of 3 meV Śliwa and Dietl (2023) shows the range of Fermi level energies for which the Hall resistance will assume the quantized value Rxy= Liu et al. (2008), corresponding to the Chern number C = 1. The gate voltage width of the Hall resistance plateau will depend on the density of the in-gap localized states in a given QW Dietl (2023a).
VII Conclusions
According to theoretical studies summarized in the present and the companion paper Śliwa and Dietl (2023), the predicted magnitude and often sign of magnetic coupling between cation-substitutional Cr ions in II-VI compounds depends on the approach employed. In particular, our results show that differing conclusions about the sign of the interaction (FM vs. AFM) come out from the GGA+ approach and the hybrid functional method. This fact reflects a delicate balance between FM superexchange and mostly AFM Bloembergen-Rowland and two-electron terms Śliwa and Dietl (2023). Altogether, the data point out that while coupling between Mn ions is AFM, the FM interactions take over in the V case, but Cr in II-VI constitutes a borderline situation – FM and AFM components are of a similar strength. This may mean that the observation of the QAHE may require the application of a magnetic field to polarize Cr spins. At the same time, it might be possible to manipulate with magnetism and, thus, with topological properties by electric field, strain, pressure, and chemical content.
Another relevant question concerns the position of states brought about by magnetic impurities, which determines their electrical activity and the degree of hybridization with the band states. It has been known for a long time Dietl and Ohno (2014) that, in II-VI compounds, Mn donor and acceptor states reside deep in the valence band and high in the conduction band, respectively, so that Mn acts as an isoelectronic impurity, the fact reconfirmed by our data. Because the - interaction between Te orbitals and Hg levels, the top of the valence is high in HgTe, which results in a relatively small value of the - exchange integral in HgTe compared to other II-VI compounds Dietl and Ohno (2014); Autieri et al. (2021).
Furthermore, our results confirm the experimental finding that the V and Cr donor states are the gap of CdTe. However, according to the outcome of our computations, V states lie the conduction band of HgTe, meaning that V ions act as resonant donors. If this were the case, it would be hard to shift the Fermi level to the topological gap in (Hg,V)Te QWs. Our results indicated that the Cr donor states reside in the HgTe valence band, so that Cr ions constitute isoelectronic impurities. Since, however, those states are relatively close to the top of the valence band, we predict a relatively sizable magnitude of in (Hg,Cr)Te. According to previous Liu et al. (2008) and recent works Śliwa and Dietl (2023), a large ratio of to the - exchange integral in HgTe/(Cd,Hg)Te QWs doped with transition metals either in the QW or barriers, results in such an ordering of spin subbands, which impedes the appearance of the QAHE. Fortunately, if magnetization is tilted out of the growth direction, the spin-orbit interaction reduces the - spitting of QW states without affecting the - exchange. The presence of the QAHE is predicted for the [111] direction of magnetization in the (Hg,Cr)Te QWs Śliwa and Dietl (2023). In any case, a close energetic proximity of states and the Fermi level opens doors for a new physics in both (Hg,V)Te and (Hg,Cr)Te, not encountered in the case of (Hg,Mn)Te. A pool of other possibilities are offered by engineering superlattices or heterostructures involving TM-doped Hg-based topological systems, some of possibilities being indicated elsewhere Islam et al. (2022, 2023).
@@ Acknowledgments
This research was supported by the Foundation for Polish Science through the IRA Programme co-financed by the EU within SG OP and EFSE (project “MagTop” no. FENG.02.01-IP.05-0028/23), by funds from the state budget allocated by the Minister of Science (Poland) as part of the Polish Metrology II programme project no. PM-II/SP/0012/2024/02. We acknowledge the access to the computing facilities of the Interdisciplinary Center of Modeling at the University of Warsaw, Grant G84-0, GB84-1, and GB84-7. We acknowledge the CINECA award under the ISCRA initiative IsC85 ”TOPMOST” and IsC93 ”RATIO” grant, for the availability of high-performance computing resources and support. We acknowledge the access to the computing facilities of the Poznan Supercomputing and Networking Center Grant No. 609.
@@ Appendix A Transition metal doped compounds with 25% and 50% of the exact exchange
In this Appendix, we present band structures and the magnetic properties for TM-doped CdTe and HgTe obtained employing the hybrid functional with and of the exact exchange. The band structure results displayed in Figs. 6, 7, and 8 further substantiate the use of in the main body of the text. In particular, in the case of Hg0.875Mn0.125Te, the experimental value of the band gap eV Furdyna (1988) is close to the computational data for (Fig. 3) but is too small and too large for and 0.5, respectively, as shown in Figs. 7 and 8.
@@
| Cd1-xVxTe | -38.1 | 22.8 | 0.6 |
| Cd1-xCrxTe | -10.5* | -16.8* | -4.8 |
| Cd1-xMnxTe | -21.9 | -21.8 | -0.5 |
| Hg1-xVxTe | 47.6 | 51.5 | -1.6 |
| Hg1-xCrxTe | -77.9 | -35.9 | -5.1 |
| Hg1-xMnxTe | -22.3 | -22.2 | -2.0 |
@@
| Cd1-xCrxTe NO JT | -218.7 | -160.3 |
| Cd1-xCrxTe JT/2 | -62.3 | 24.1 |
| Hg1-xCrxTe NO JT | 74.0 | 7.8 |
| Hg1-xCrxTe JT/2 | 88.6 | -22.7 |
Appendix B Magnetic configurations
In this Section, for completeness, we report the graphical representation of the magnetic configurations investigated in the paper. We have used 222 supercells with two magnetic impurities and we have investigated six different configurations, three in the ferromagnetic case and three in the antiferromagnetic case. The FM configurations are reported in Fig. 9, while the AFM configurations are shown in Fig. 10. In the unit cell, we have 8 cations and 8 anions. In the first configuration, the two magnetic dopants replace two cations first-neighbor, in the second configuration the magnetic dopants are second-neighbor, while in the third one, the magnetic dopants are third-neighbors. Since there are also the atoms in the repeated unit cell to be considered, (i=1,2,3) are not exactly proportional to the first-neighbor, second-neighbor and third-neighbor magnetic couplings. However, they give a strong indication on these magnetic couplings.
Appendix C Results within GGA + U approximation
In Tab. 7 we report the values obtained by using GGA + calculations. As we can see from Tab. 7, in the GGA approximation, in the case of doping with Mn, we have antiferromagnetic couplings for both HgTe and CdTe. In the case of doping with Cr and V, the couplings are ferromagnetic. The values of the third neighbor energy differences for Hg1-xCrxTe and Cd1-xCrxTe in GGA + U are unphysical. The antiferromagnetism of the couplings in the case of Mn-doped systems is still present when we use the HSE functional, therefore this property is robust, and also because already within GGA + we get the experimental insulating configuration for these systems. In GGA+ the ferromagnetism is overestimated due to the overestimation of the metallicity. In Tab. 8 we report the results obtained with GGA + U for the case of doping with Cr with JT distortions included. As in the main body of the paper, with JT/2 we indicate that we are considering the average of the positions of the atoms without and with the Jahn-Teller distortion.
| E1 | E2 | E3 | |
| Cd1-xCrxTe | 63.5 | 215.9 | 107.6 |
| Cd1-xVxTe | 497.6 | 28.2 | -1.7 |
| Cd1-xMnxTe | -15.7 | -15.6 | -0.7 |
| Hg1-xCrxTe | 168.3 | 178.4 | -36.1 |
| Hg1-xVxTe | 44.1 | 44.2 | -1.8 |
| Hg1-xMnxTe | -13.0 | -12.9 | 0.4 |
| E1 | E2 | E3 | |
| Cd1-xCrxTe NO JT | 63.5 | 215.9 | 107.6 |
| Hg1-xCrxTe NO JT | 168.3 | 178.4 | -36.1 |
| Cd1-xCrxTe JT/2 | 70.3 | 220.9 | 100.9 |
| Hg1-xCrxTe JT/2 | 171.5 | 174.4 | -32.9 |
| Cd1-xCrxTe JT | 74.8 | 140.1 | 91.4 |
| Hg1-xCrxTe JT | 171.6 | 170.5 | -30.4 |